# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v2

@@ 3.3 Application to the case study
In the motivating case, the primary goal is to assess the effectiveness of a newly introduced traffic fee aimed at regulating vehicular traffic in a specific urban area by limiting vehicle circulation. Given this scope, key variables for analysis include vehicular traffic flow measurements alongside additional metrics on vehicle emissions and air quality, which will act as performance metrics for the restrictions on circulation and on pollution. The policy is defined by control parameters that quantify the regulatory restriction across temporal boundaries, affected user bases, and economic burdens imposed by the measure. Furthermore, the model captures user reactions as a function of both behavioral attributes and the defining properties of the policy. This section presents these variables and the interconnected relationships that allow moving from one to the other, as summarized in Figure 2.
Three core variables are observed and monitored in the baseline and modified scenarios. The emissions measure the overall pollutants generated by the vehicles circulating in the area. Moreover, two other quantities control the number of vehicles and thus measure the effects of policy limitations: the inflow represents the number of vehicles entering the area affected by the policy regulation, and traffic defines the number of vehicles circulating in the area. These indices can be derived from various real-world mobility data sources, including traffic counting sensors that monitor vehicle flows, origin-destination matrices that capture travel patterns between different zones, and GPS or mobile phone data that provide movement trajectories. Their elaboration leads to the characterization of the current baseline scenario.
@@
The policy enacts an area-based traffic regulation for which the vehicular circulation in a designated area is limited. In this example, the regulation imposes a fee on all vehicles circulating into the area during the daily hours. The cost depends on the type of vehicle in circulation: to achieve emission reduction goals alongside traffic reduction, different penalty levels can be applied to vehicles based on their emission standards, for instance, with higher costs assigned to older, more polluting classes. Moreover, the policy provides payment exemptions for vulnerable social groups.
@@
In response to the implementation of the regulatory policy, affected travelers reassess their transportation habits and may modify them accordingly. The model introduces behavioral parameters, designed to define users’ attitudes and preferences and account for the most likely behavior strategies. Among them, the most plausible, included in the model, are: rigidity (i.e., portion of travelers, among the affected ones, that continues with their usual mobility habits), time-shifting (i.e., fraction of affected travelers that accepts shifting their trip to an earlier time before the policy becomes active or to a later time after the policy ends), and mode-shifting (i.e., portion of travelers that decides to change the mode of transport from private vehicles towards public transport means). If none of the other alternatives are viable, travelers may give up their usual trip and become lost vehicles. This case models trip cancellations, for instance, but also changes to other modes of transport not monitored in this model (e.g., foot, bicycle).
@@
The application of the policy and its effects on the user travel behaviours initiate a sequential process that progressively alters inflow, traffic, and emission levels in the modified scenario. The effects on the inflow follow directly from the policy implementation. In the new scenario, the inflow varies from the base one with time: it may be reduced when the policy is applied, and there remain only exempted and rigid travelers; on the other hand, it may be increased in the hours of policy inactivation, when anticipating and postponing travelers are to be added to the base inflow. When aggregating traffic from all vehicles, both those starting their journey inside the area and those arriving from outside, the total traffic volume changes in accordance with the variations in the number of vehicles. The new scenario is hence associated with a modified value of total traffic in the area per time instant. This causes changes in the vehicular emissions from the area, which vary according to the traffic. The policy and the subsequent scenario modification affect other measurable quantities, which could be important for a complete policy assessment. For instance, within the transport domain, the extent of the modal shift towards public transport can be assessed.
@@
Figure 2: Conceptual diagram of the case study, showing the main quantities of interest (i.e., inflow, traffic, and emissions) in the current and modified scenarios. Blue squares represent modeling steps that map inputs to outputs, allowing for the assessment of the modified quantities
@@ 4.1 Scenario definition
The analysis focuses on an urban area, of which a portion of interest is designated for the policy implementation, and other external areas surround it. The analysis considers one reference working day, with measures taken during the 24 hours of the day with a sampling frequency $\nu$. The policy is assumed to be applied on the reference day during designated time slots, with a starting time $t_{s}$ and an ending time $t_{e}$. A fraction of vehicles is exempted from its action, $F_{e}$, while it targets non-exempt vehicles with the fee amount varying according to the vehicle emission standard, $\{C_{0l}\}_{l=0,\dots,n_{l}}$.
Among the quantities of interest, $I(t)$ represents the vector of inflows, that is, the number of vehicles entering the area in each time interval $t$. In the current base scenario, these inflow values are treated as known parameters, obtained through estimation procedures applied to real traffic data, while in the modified scenario, their values are estimated with the application of the behavioral models explained in Section 4.2. The other two core variables are $T(t)$, representing the number of circulating vehicles in the area during time interval $t$, and $E(t)$, which denotes the emissions produced by all vehicles circulating within the area during time interval $t$. The estimation of these quantities is presented in Section 4.3 and Section 4.4, respectively, and follows from the known inflow patterns combined with the emission level distribution of the active vehicle fleet.
@@ 4.2 Inflow modifications
The affected travelers select among these strategies reflecting the trade-offs that they make between convenience, cost, and compliance with the policy. In general, the likelihood of each behavioral option is evaluated by looking at the difference between the effort required by the policy and the effort that travelers are willing to accept. “Effort” here refers to the general cost or burden imposed by the policy rather than solely monetary expense. For instance, those considering advancing their departure time will weigh the burden of schedule changes against their willingness to travel earlier; those evaluating modal shift will assess their effort to favour alternative transport over private vehicles; while those contemplating unchanged behavior will compare the monetary cost imposed by the policy against their willingness to pay. The probability of each response is modeled through distinct formulations, since the nature of the decision process varies across the different behavioral options. Then, the optimal behavior is selected either as the single option with minimal effort or as one among several options that each require an acceptably small effort. The following paragraphs discuss behavioral strategies and choice strategies, modeling both the propensity for each behavioral option and the final selection of a strategy.
@@
Let $X$ denote the variable of acceptable effort, that is, the maximum effort that individual travelers are willing to admit. Let us assume that this variable follows an exponential distribution of density $f(x)=\lambda\exp\{-\lambda x\}$. Among the behavioral parameters of the model, it is possible to specify the median of this distribution, $X_{50\%}$, that corresponds to the acceptable effort for 50% of travelers, and that is related to the parameter $\lambda$ through the formula $\lambda=\ln(2)/X_{50\%}$.
@@
Taking this distribution into account, once the effort required by the access policy $X_{0}$ is known, it is possible to compute the probability that the required effort is less than the acceptable effort. This probability also represents the likelihood that such effort will be accepted by travelers, thereby identifying the proportion of travelers ready to adopt the behavioral strategy.
@@
$\mathbb{P}[\text{accept the strategy}]=\mathbb{P}[X\geq X_{0}]=\exp\left\{-\frac{\ln(2)}{X_{50\%}}\cdot X_{0}\right\}$ (1)
@@
Equation 1 is applied to estimate the probability that travelers will maintain their travel routine and demonstrate travel rigidity. In this context, the effort corresponds to the monetary cost of the policy per emission level $l$, $C_{0l}$, compared with the maximum acceptable cost, $C$, exponentially distributed with median $C_{50\%}$. Using Equation 2, it is possible to calculate the probability of accepting rigidity per emission level as $\mathbb{P}[C\geq C_{0l}]$ during the policy application.
@@
$\mathbb{P}[\text{rigidity for Euro-level }l]=\mathbb{P}[C\geq C_{0l}]=\exp\left\{-\frac{\ln(2)}{C_{50\%}}\cdot C_{0l}\right\}$ (2)
@@
By reweighting the probability with the proportion of vehicles belonging to each emission level class, the probability of overall rigidity acceptance is found, as in Equation 3.
$\mathbb{P}[\text{rigidity}]=\sum_{l}P_{l}\cdot\mathbb{P}[\text{rigidity for Euro-level }l]=\sum_{l}P_{l}\cdot\mathbb{P}[C\geq C_{0l}]$ (3)
@@
The postponement of trips is evaluated by assessing the effort of a time variation from the current travel starting time to the policy ending time, $\Delta_{te}(t)=t_{e}-t$ $\forall t\in[t_{s},t_{e}]$. The time variation is to be compared with the maximum acceptable time variation, $\Delta_{p}$, exponentially distributed with median $\Delta_{p50\%}$. As $\Delta_{te}$ depends on the time, the probability of postponement is also time-dependent on each starting time interval $t$ throughout Equation 4.
@@
$\mathbb{P}[\text{postponement in time }t]=\mathbb{P}[\Delta_{p}\geq\Delta_{te}(t)]=\exp\left\{-\frac{\ln(2)}{\Delta_{p50\%}}\cdot\Delta_{te}(t)\right\}$ (4)
@@
Similarly, Equation 5 uses the acceptance anticipation $\Delta_{a}$, the required anticipation per time interval $\Delta_{fs}(t)$, and the median parameter $\Delta_{a50\%}$, to compute the probability of accepting anticipation. However, to fully avoid the circulation policy, the trip should be anticipated to be completed before the policy starts. Consider the travel time in the area, modeled as a random variable with a geometric distribution $\tau\sim\mathcal{G}eom(p=\tfrac{1}{\bar{\tau}})$ over the support $\mathbb{N}_{>0}$. Then the probability of anticipation can be found as in Equation 5.
@@
where $\varepsilon=p\,e^{-{\ln(2)}/{\Delta_{a50\%}}}[1-(1-p)e^{-{\ln(2)}/{\Delta_{a50\%}}}]^{-1}$.
@@
The probability of the change in the mode of transportation is modeled through a logit model (Ben-Akiva and Bierlaire, 1999). It is based on the principle that the probability of changing between two modes of transport depends on the relative effort of the options. In this case, since focusing on the shift from private to public transport, the model incorporates covariates related to the availability of the public transport system, as well as indicators of the overall convenience of making the change, to assess the modal shift proportion. As the public transport distribution is not homogeneous in the city, we characterized subzones of the metropolitan areas with their indicators of availability and convenience, and applied independent modal shift formulations for each travel zone. Consequently, each external origin zone generating trips to the area is assigned a specific probability of modal shift.
Let $p_{z}$ be the probability of accepting a mode shift to the public transport when starting the trip in zone $z$. Then, its logit is related to the linear combination of the coefficients $\beta_{j}$ and the transport convenience covariates $x_{zj}$.
$\ln\frac{p_{z}}{1-p_{z}}=\beta_{0}+\beta_{1}x_{z1}+\beta_{2}x_{z2}+\beta_{3}x_{z3}+\beta_{4}x_{z4}$ (6)
Two covariates, $x_{z1}$ and $x_{z2}$, represent the objective characteristics of public transport supply and their adequacy relative to travelers’ needs per zone: frequency and capillarity. The frequency is defined as the passage frequency of public transport at stops, and is considered an index of service availability. For each stop and line passing through it, the average number of hourly passages is calculated. By then aggregating this value by zone, the average passage frequency of the vehicles at the stops in the area is obtained. The capillarity is defined as the proportion of the transportation network covered by the public transport system, and is considered a proxy of the service accessibility. It is calculated per zone as the ratio of the transport network length served by public transit routes, determined by matching the network to the public transit routes, to the total transport network length.
Two additional indices, $x_{z3}$ and $x_{z4}$, evaluate the convenience of public transport compared to private transport: cost and time difference. The cost of the public transportation service is the fare that a user would pay to travel via public transportation from any zone to the area of interest, evaluating the economic convenience of the mode-shift. The difference in travel times when using public transport and when using private means, then, assesses the temporal convenience of the mode-shift.
@@
For each zone of origin, the probability of the modal shift is given by Equation 7.
@@
$\mathbb{P}[\text{modal shift from zone }z]=\text{logit}^{-1}\{\beta_{0}+\beta_{1}x_{z1}+\beta_{2}x_{z2}+\beta_{3}x_{z3}+\beta_{4}x_{z4}\}$ (7)
@@
By reweighting the probability with the proportion of vehicles starting in each zone, $Z_{z}$, the probability of overall modal shift acceptance is found as in Equation 8.
$\mathbb{P}[\text{modal shift}]=\sum_{z}Z_{z}\cdot\mathbb{P}[\text{modal shift from zone }z]$ (8)
Optimal strategy choice:
After evaluating the available options, travelers decide what to choose, and the model accounts for the different evaluation methods they use. For instance, Travelers may assess options either sequentially, considering them one after the other and giving priority to certain behaviors, or simultaneously, comparing all alternatives to identify the most suitable choice. In this paper, travelers are assumed to consider the options simultaneously and independently of one another, and then decide their strategy as the one that they prefer. This represents a modeling choice; however, the framework can readily accommodate alternative decision-making models, such as the sequential evaluation strategy or hybrid approaches that combine multiple decision rules.
The Equations 3, 5, 4, and 8 report the marginal probabilities that a specific strategy is acceptable for travelers. Since the acceptability of a strategy is independent of the others, the fact that a traveler finds option $X$ acceptable does not influence the probability that they also find strategy $Y$ acceptable. As such, a traveler can find themselves in different situations: they may find acceptable a single strategy, or no strategy, or any combination of strategies, without one excluding the others in terms of acceptability level. In the case of acceptance of multiple strategies, the traveler must nonetheless choose which one to concretely implement. The model assumes that in this situation, the traveler is indifferent among the acceptable options and selects one at random with uniform probability. For example, if $k$ acceptable solutions, each has a probability $1/k$ of being chosen.
@@
Let’s take an example with the calculation of the fraction of rigid travelers. This will be equal to the fraction of non-exempt individuals who accept rigidity but do not accept any other strategy, plus the portion of non-exempt individuals who accept both rigidity and another strategy and then randomly choose the former with probability $1/2$. To this, we add the portion of non-exempt individuals with three accepted strategies, including rigidity, who randomly choose rigidity over the other solutions with probability $1/3$. Finally, we include cases where non-exempt individuals find all four strategies acceptable and, among them, choose rigidity with probability $1/4$. Note that, since the marginal probabilities of anticipation and postponement are time-dependent, also the fraction of rigid varies with time $t\in[t_{s},t_{e}]$. In general, given any strategy $X$ chosen from those available, $E=\{\text{rigidity},\text{anticipation},\text{postponement},\text{mode-shift}\}$, the fraction of travelers choosing such a strategy is calculated as in Equation 9.
$F_{X}(t)=(1-F_{e})\cdot\sum_{S\subseteq E;X\in S}\frac{1}{|S|}\prod_{Y\in S}\mathbb{P}[Y](t)\prod_{Z\notin S}(1-\mathbb{P}[Z](t))$ (9)
@@
Those that do not accept any of the strategies become part of the lost vehicles, either canceling the trip, or changing to other transport modes rather than public transport, or choosing other options not accounted for in the model. The fraction of lost vehicles is computed as the fraction of non-exempted vehicles, multiplied by the probability of not accepting any of the behavioral strategies proposed, as in Equation 10.
@@
Having determined the proportion of vehicles adopting each behavioral strategy, most categories can be directly processed: rigid and exempt vehicles contribute to the calculation of the inflow during policy time slots, while mode-shifted and lost vehicles are removed from the analysis. On the other hand, time-shifting vehicles require an additional layer of analysis. Besides deciding whether to shift their departure time, these drivers must also determine how much to anticipate or delay their trips. For instance, a driver who chooses to avoid a policy by departing earlier faces the second decision of how far in advance they should actually travel. This temporal reallocation introduces another modeling phase, as these vehicles must remain in the daily total count while being redistributed across different time intervals according to specific shifting behaviors.
This model accounts for a time-shifting strategy that separates the shifting decision into two stages. First, travelers decide whether to accept time-shifting based solely on avoiding the policy period, regardless of the specific alternative timing. Then, those who accept are redistributed to new departure times before or after the policy period without considering their original departure preferences.
Specifically, from the first phase, the fraction and, therefore, the number of vehicles that choose a time shift are identified. Then, the total of vehicles anticipating $TN_{a}$ and postponing $TN_{p}$ are calculated as in Equation 11.
@@
Through the second phase, the redistribution of these totals is modeled both before and after the policy’s duration. The reasoning applies to both anticipation and postponement, but let us now consider the case of postponement. The random variable $\Delta_{pr}$ indicates the time variation that travelers are willing to make when postponing their redistribution, specifically, the time span between their new trip’s starting time and the policy’s end. The exponential distribution is chosen for $\Delta_{pr}$ because it naturally reflects decreasing willingness as time shifts increase (most travelers accept small schedule changes, but fewer accept larger disruptions). It has median $\Delta_{pr50\%}$, scale parameter equal to $\lambda=\frac{\ln(2)}{\Delta_{pr50\%}}$, and the density $f_{pr}(x)=\lambda\exp\{-\lambda x\}$. When determining how many trips will be moved to a specific time interval $t$, requiring a time shift from the policy end $\Delta_{fe}(t)$, the analysis requires the density of travelers whose personal tolerance exactly matches this required shift. This density directly gives the rate at which trips concentrate at time interval $t$, computed as in Equation 12. The same reasoning applies symmetrically to anticipations, where we consider the random variable $\Delta_{ar}$ with median $\Delta_{ar50\%}$, density $f_{ar}$, and the time variation is measured until the start of the policy period, denoted as $\Delta_{ts}(t)$.

### v3

@@ 3.3 Application to the case study
In the motivating case, the primary goal is to assess the effectiveness of a newly introduced regulations concerning vehicular traffic in a specific urban area. Even if the ultimate goal is to build a general framework for traffic policy appraisal, in order to keep the discussion focues, the case is limited a regulation concerning the introduction of a traffic fee meant to limit vehicle circulation. Given this scope, key variables for analysis include vehicular traffic flow measurements alongside additional metrics on vehicle emissions and air quality, which will act as performance metrics for the restrictions on circulation and on pollution. The policy is defined by control parameters that quantify the regulatory restriction across temporal boundaries, affected user bases, and economic burdens imposed by the measure. Furthermore, the model captures user reactions as a function of both behavioral attributes and the defining properties of the policy. This Section presents these variables and the interconnected relationships that allow moving from one to the other, as summarized in Figure 2.
Three core variables are observed and monitored in the as-is and what-if scenarios. Two quantities control the number of vehicles and thus measure the direct effects of policy limitations: the inflow represents the number of vehicles entering the area affected by the policy regulation, and traffic defines the number of vehicles circulating in the area. Then, the emissions measure the overall pollutants generated by the vehicles circulating in the area, sheding lights on the policy impact on health and environment.
@@
The policy enacts an area-based traffic regulation for which the vehicular circulation in a designated area is limited. In this example, the regulation imposes a fee on all vehicles circulating into the area during the daily hours. The cost depends on the type of vehicle in circulation: to achieve emission reduction goals alongside traffic reduction, different penalty levels can be applied to vehicles based on their emission standards, for instance, with higher costs assigned to more polluting classes. Moreover, the policy provides payment exemptions for vulnerable social groups.
@@
In response to the implementation of the regulatory policy, affected travelers reassess their transportation habits and may modify them accordingly. The model introduces behavioral parameters, designed to define users’ attitudes and preferences and account for the most likely behavior strategies. Among them, the most plausible, included in the model, are: rigidity (i.e., continuing with the usual mobility habits), time-shifting (i.e., shifting the trip to an earlier time before the policy becomes active or to a later time after the policy ends), and mode-shifting (i.e., changing the mode of transport from private vehicles towards public transport means). If none of the alternatives are viable, the model incorporates the lost strategy; this represents trip cancellations, for instance, but also changes to other modes of transport not monitored in this model (e.g., foot, bicycle).
@@
The application of the policy and its effects on the user travel behaviors initiate a sequential process that progressively alters inflow, traffic, and emission levels in the what-if scenario. The effects on the inflow follow directly from the policy implementation. In the new scenario, the inflow varies from the base one with time: it may be reduced when the policy is applied and there remain only exempted and rigid travelers; on the other hand, it may be increased in the hours of policy inactivation, when anticipating and postponing travelers are to be added to the base inflow. When aggregating traffic from all vehicles, both those starting their journey inside the area and those arriving from outside, the total traffic volume changes in accordance with the variations in the number of vehicles. The new scenario is hence associated with a modified value of total traffic in the area per time instant. This causes changes in the vehicular emissions from the area, which vary according to the traffic. The policy and the subsequent scenario modification affect several key metrics necessary for a full policy evaluation. Illustrative examples include direct transport metrics, such as the extent of modal shift, and various indirect impacts, such as social mobility constraints and broader economic consequences for tourism and commerce. While not the primary focus of this study, these secondary effects are an important dimension of the phenomenon and are currently being examined in ongoing work, with further analyses planned in future research.
TrafficInflowEmissionsInflowTrafficEmissionsControl ParametersBehavioral ParametersAs-is ScenarioWhat-if Scenario other indirect impacts other indirect impacts other indirect impacts
@@
Figure 2: Conceptual diagram of the case study, showing the main quantities of interest (i.e., inflow, traffic, and emissions) in the as-id and what-if scenarios. Blue squares represent modeling steps that map inputs to outputs, allowing for the assessment of the modified quantities.
@@ 4.1 Scenario definition
The analysis focuses on an urban area, of which a portion of interest is designated for the policy implementation, and other external areas surround it. The analysis considers one reference working day, with measures taken during the 24 hours of the day with a sampling frequency $\nu$ (e.g., every 5 minutes). The policy is assumed to be applied on the reference day during designated time slots, with a starting time $t_{s}$ and an ending time $t_{e}$. A fraction of vehicles is exempted from its action, $F_{e}$, while it targets non-exempt vehicles with the fee amount varying according to the vehicle emission standard: assuming there are $n_{l}$ standards, the costs are $\{C_{0l}\}_{l=0,\dots,n_{l}}$.
Among the quantities of interest, $I(t)$ represents the vector of inflows, that is, the number of vehicles entering the area in each time interval $t$. In the as-is base scenario, these inflow values are treated as known parameters, obtained through estimation procedures applied to real traffic data, while in the what-if scenario, their values are estimated with the application of the behavioral models explained in Section 4.2. The other two core variables are $T(t)$, representing the number of circulating vehicles in the area during time interval $t$, and $E(t)$, which denotes the emissions produced by all vehicles circulating within the area during time interval $t$. The estimation of these quantities is presented in Section 4.3 and Section 4.4, respectively, and follows from the known inflow patterns combined with the emission level distribution of the active vehicle fleet. We remark that the explicit models introduced below provide one possible implementation of the framework. Other specifications can be used within the same framework, which accomodates a wide range of alternative modeling assumptions.
@@ 4.2 Inflow modifications
The evaluation of these strategies reflects the trade-offs between convenience, cost, and compliance with the policy. In general, the likelihood of each behavioral option is evaluated by looking at the difference between the effort required by the policy and the maximum acceptable effort. “Effort” here refers to the general cost or burden imposed by the policy rather than solely monetary expense. For instance, advancing the departure time will weigh the burden of schedule changes against the willingness to travel earlier; evaluating modal shift will assess the effort to favour alternative transport over private vehicles; while contemplating unchanged behavior will compare the monetary cost imposed by the policy against the willingness to pay.
The probability of each behavioral response is modeled using distinct formulations, reflecting the aggregate behavioral tendencies observed in similar contexts. Rather than representing individual-level decision processes, these formulations capture the emergent collective behavior resulting from heterogeneous preferences and constraints. Then, the optimal behavior is identified based on the acceptability likelihood of the strategies. The following paragraphs discuss the behavioral strategies and the corresponding modeling framework adopted to represent both the propensity toward each option and the resulting aggregate selection patterns.
@@
Let $X$ denote the variable of maximum acceptable effort. Let us assume that this variable follows an exponential distribution of density $f(x)=\lambda\exp\{-\lambda x\}$. Among the behavioral parameters of the model, it is possible to specify the median of this distribution, $X_{50\%}$, that corresponds to the acceptable effort for 50% of travelers, and that is related to the parameter $\lambda$ through the formula $\lambda=\ln(2)/X_{50\%}$.
@@
Then, once the effort required by the access policy $X_{0}$ is known, it is possible to compute the probability that the required effort is less than the acceptable effort, as in Equation 1. This probability also represents the acceptance rate of the strategy, thereby defining the share of travelers allocated to that behavioral response.
@@
$\mathbb{P}[\text{accept the strategy}]=\mathbb{P}[X\geq X_{0}]=\exp\left\{-\frac{\ln(2)}{X_{50\%}}\cdot X_{0}\right\}.$ (1)
@@
Equation 1 is applied to estimate the probability of accepting rigidity. In this context, the effort corresponds to the monetary cost of the policy per emission level $l$, $C_{0l}$, compared with the maximum acceptable cost, $C$, exponentially distributed with median $C_{50\%}$. Using Equation 2, it is possible to calculate the probability of accepting rigidity per emission level as $\mathbb{P}[C\geq C_{0l}]$ during the policy application:
@@
$\mathbb{P}[\text{rigidity for Euro-level }l]=\mathbb{P}[C\geq C_{0l}]=\exp\left\{-\frac{\ln(2)}{C_{50\%}}\cdot C_{0l}\right\}.$ (2)
@@
By reweighting the probability with the proportion of vehicles belonging to each emission level class, the probability of overall rigidity acceptance is found, as in Equation 3:
$\mathbb{P}[\text{rigidity}]=\sum_{l}P_{l}\cdot\mathbb{P}[\text{rigidity for Euro-level }l]=\sum_{l}P_{l}\cdot\mathbb{P}[C\geq C_{0l}].$ (3)
@@
The postponement of trips is evaluated by assessing the effort of a time variation from the current travel starting time to the policy ending time, $\Delta_{te}(t)=t_{e}-t$ $\forall t\in[t_{s},t_{e}]$. The time variation is to be compared with the maximum acceptable time variation, $\Delta_{p}$, exponentially distributed with median $\Delta_{p50\%}$. As $\Delta_{te}$ depends on the time, the probability of postponement is also time-dependent on each starting time interval $t$ throughout Equation 4:
@@
$\mathbb{P}[\text{postponement in time }t]=\mathbb{P}[\Delta_{p}\geq\Delta_{te}(t)]=\exp\left\{-\frac{\ln(2)}{\Delta_{p50\%}}\cdot\Delta_{te}(t)\right\}.$ (4)
@@
Similarly, Equation 5 uses the acceptance anticipation $\Delta_{a}$, the required anticipation per time interval $\Delta_{fs}(t)$, and the median parameter $\Delta_{a50\%}$, to compute the probability of accepting anticipation. However, to fully avoid the circulation policy, the trip should be anticipated to be completed before the policy starts. Consider the travel time in the area, modeled as a random variable with a geometric distribution $\tau\sim\mathcal{G}eom(p=\tfrac{1}{\bar{\tau}})$ over the support $\mathbb{N}_{>0}$. Then the probability of anticipation can be found as in Equation 5:
@@
where $\varepsilon=p\,\exp\{-{\ln(2)}/{\Delta_{a50\%}}\}[1-(1-p)\exp\{-{\ln(2)}/{\Delta_{a50\%}}\}]^{-1}$.
@@
The probability of the change in the mode of transportation is modeled through a logit model (Ben-Akiva and Bierlaire, 1999). It is based on the principle that the probability of changing between two modes of transport depends on the relative effort of the options. In this case, since focusing on the shift from private to public transport, the model incorporates covariates related to the availability of the public transport system (PTS), as well as indicators of the overall convenience of making the change, to assess the modal shift proportion. As the public transport distribution is not homogeneous in the city, we characterize subzones of the metropolitan areas with their indicators of availability and convenience, and apply independent modal shift formulations for each travel zone. Consequently, each external origin zone generating trips to the area is assigned a specific probability of modal shift.
Among the considered covariates, two of them represent the objective characteristics of public transport supply and their adequacy relative to travelers’ needs per zone: $x_{1z}$ is the frequency and $x_{2z}$ is the capillarity in zone $z$. The frequency is defined as the passage frequency of public transport at stops, and is considered an index of service availability. For each stop and line passing through it, the average number of hourly passages is calculated. By then aggregating this value by zone, the average passage frequency of the vehicles at the stops in the area is obtained. The capillarity is defined as the proportion of the transportation network covered by the PTS, and is considered a proxy of the service accessibility. It is calculated per zone as the ratio of the transport network length served by public transit routes, determined by matching the network to the public transit routes, to the total transport network length.
Two additional covariates, then, evaluate the convenience of public transport compared to private transport: $x_{3z}$ is the cost and $x_{4z}$ is the time difference in the zone $z$. The cost of the public transportation service is the fare that a user would pay to travel via public transportation from any zone to the area of interest, evaluating the economic convenience of the mode-shift. The difference in travel times when using public transport and when using private means, then, assesses the temporal convenience of the mode-shift.
Let $p_{z}$ be the probability of accepting a mode shift to the public transport when starting the trip in zone $z$. Then, its logit is related to the linear combination of the transport covariates, $x_{1z},\dots,x_{4z}$., as in the following formulation:
$\text{logit}(p_{z})=\ln\frac{p_{z}}{1-p_{z}}=\beta_{0}+\sum_{i=1}^{4}\beta_{i}x_{iz}.$ (6)
@@
For each zone of origin, the probability of the modal shift is given by Equation 7:
@@
$\mathbb{P}[\text{modal shift from zone }z]=\text{logit}^{-1}\left(\beta_{0}+\sum_{i=1}^{4}\beta_{i}x_{iz}\right).$ (7)
@@
By reweighting the probability with the proportion of vehicles starting in each zone, $Z_{z}$, the probability of overall modal shift acceptance is found as in the following Equation:
$\mathbb{P}[\text{modal shift}]=\sum_{z}Z_{z}\cdot\mathbb{P}[\text{modal shift from zone }z].$ (8)
Aggregate strategy-choice pattern:
After defining the set of available behavioral options, the model represents how these alternatives are comparatively evaluated at the aggregate level. Rather than describing individual cognitive processes, the framework specifies alternative evaluation rules that approximate collective behavioral patterns observed in similar contexts. For example, options may be assessed either sequentially (by assigning priority to certain behaviors) or simultaneously (by comparing all alternatives within a common evaluative structure). In this study, a simultaneous and independent evaluation of options is adopted as a modeling assumption, and the selected strategy corresponds to the alternative associated with the lowest perceived effort at the aggregate level. This represents a modeling choice; however, the framework can readily accommodate alternative models, such as the sequential evaluation strategy or hybrid approaches that combine multiple decision rules.
Equations 3, 5, 4, and 8 report the marginal probabilities that a specific strategy is acceptable within the modeling framework. These probabilities are specified independently, such that the acceptability of one strategy does not affect the acceptability of another. As such, the framework allows for multiple possible configurations: a single strategy may be acceptable, none may be acceptable, or several strategies may simultaneously satisfy the acceptability condition. In the case of acceptance of multiple strategies, the model requires an allocation rule to determine the implemented outcome. Specifically, each acceptable strategy is assigned a weight that determines its probability of being selected. These weights can be uniform, reflecting a situation in which all acceptable strategies are equally likely to be chosen (e.g., analogous to a random selection), or they can reflect relative probabilities that capture preferences, likelihoods, or other decision criteria.
@@
Note that, since the marginal probabilities of anticipation and postponement are time-dependent, also the fraction of rigid varies with time $t\in[t_{s},t_{e}]$. In general, given any strategy $X$ chosen from those available, $E=\{\text{rigidity},\text{anticipation},\text{postponement},\text{mode-shift}\}$, the fraction of travelers with such a strategy is calculated in Equation 9, obtaining respectively $F_{rig},F_{ant},F_{post},$ and $F_{mode}$:
$F_{X}(t)=(1-F_{e})\cdot\sum_{S\subseteq E;X\in S}w(X,S,t)\prod_{Y\in S}\mathbb{P}[Y](t)\prod_{Z\notin S}(1-\mathbb{P}[Z](t)),$ (9)
where $S\subseteq E$ is any combination of strategies including $X$; and $w(X,S,t)$ represents the weight associated to the acceptability of strategy $X$ within the set $S$ at time $t$. In the current implementation, we use the uniform weights: $w(X,S,t)=1/|S|$. However, as discussed, alternative weighting schemes can be used within the same framework, such as weights proportional to the marginal probabilities: $w(X,S,t)={\mathbb{P}[X](t)}/{\sum_{S}\mathbb{P}[Y](t)}$.
@@
Those for which none of the strategies are acceptable by the model become part of the lost vehicles, either canceling the trip, or changing to other transport modes rather than public transport, or choosing other options not accounted for in the model. The fraction of lost vehicles is computed as the fraction of non-exempted vehicles, multiplied by the probability that a vehicle falls into the category of non-acceptable strategies, as in Equation 10.
@@
Having determined the proportion of vehicles adopting each behavioral strategy, most categories can be directly processed: rigid and exempt vehicles contribute to the calculation of the inflow during policy time slots, while mode-shifted and lost vehicles are removed from the analysis. On the other hand, time-shifting vehicles require an additional layer of analysis. Besides representing whether vehicles shift their departure time, the model also specifies the magnitude of the anticipation or postponement. For example, for vehicles assigned to an earlier departure strategy, the model determines the corresponding timing adjustment according to a prescribed allocation rule. This temporal reallocation introduces another modeling phase, as these vehicles must remain in the daily total count while being redistributed across different time intervals according to specific shifting behaviors.
This model accounts for a time-shifting approach that separates the shifting decision into two stages. First, vehicles are classified according to whether the time-shifting strategy applies, based on the model’s criterion of avoiding the policy period, regardless of specific alternative timings. Vehicles assigned to this strategy are then redistributed to new departure times before or after the policy period, independently of their original departure times.
Specifically, using Equation 9 from the first phase, the fraction and, therefore, the number of vehicles assigned to a time shift are identified. Then, the total of vehicles anticipating $TN_{ant}$ and postponing $TN_{post}$ are calculated as in Equation 11.
@@
Through the second phase, the redistribution of these totals is modeled both before and after the policy’s duration. The reasoning applies to both anticipation and postponement, but let us now consider the case of postponement. The random variable $\Delta_{pr}$ indicates the time variation associated with the postponement strategy, specifically, the interval between the assigned new departure time and the end of the policy period. The exponential distribution is chosen for $\Delta_{pr}$, with median $\Delta_{pr50\%}$, scale parameter equal to $\lambda=\ln(2)/\Delta_{pr50\%}$, and the density $f_{pr}(x)=\lambda\exp\{-\lambda x\}$. When determining how many trips will be moved to a specific time interval $t$, the analysis requires the density corresponding to the specified timing adjustment, representing the fraction of vehicles assigned by the model to a shift from the policy end of exactly the magnitude $\Delta_{fe}(t)$. This density directly gives the rate at which trips concentrate at time interval $t$, computed as in Equation 12. The same reasoning applies symmetrically to anticipations, where we consider the random variable $\Delta_{ar}$ with median $\Delta_{ar50\%}$, density $f_{ar}$, and the time variation is measured until the start of the policy period, denoted as $\Delta_{ts}(t)$.